OpenStudy (anonymous):

I need some help with this problem. I am supposed to find the difference quotient for the function. f(x) = 3x + x^2

OpenStudy (anonymous):

I have this formula f(x + h) - f(x) over h. So I'm not sure how I'm supposed to set this up?

OpenStudy (anonymous):

You already have the function, so you need to find f(x+h). To me it looks like it might be\[f(x) = 3(x+h) + (x+h)^2=3x+3h+x^2+2hx+h^2\]\[=x^2+(3+2h)x+3h+h^2\]

OpenStudy (anonymous):

Error in entry above. The evaluation is for f(x+h), not f(x).

OpenStudy (anonymous):

Now you put it into the formula, and simplify.